Women’s heights are normally distributed with mean 63.9
inches and standard deviation 2.8 inches. Men’s heights are normally distributed with mean 68.4 inches and standard
deviation 3.0 inches. The US Navy requires that fighter pilots have heights between 62 and 78 inches.


Find the percentage of women meeting the height requirement to be a fighter pilot.
Find the percentage of men that are too short to be fighter pilots.
Answer #1

Find the percentage of women meeting the height requirement to be a fighter pilot.

Z = (X - mean)/sd = (X - 63.9)/2.8

P(62 <X < 78)

= P(-0.68 < Z < 5.04)

= 0.7513


Find the percentage of men that are too short to be fighter pilots.

Z= (Y - 68.4)/3

P(Y < 62)

= P(Z < -2.13)

= 0.0164

Please give me a thumbs-up if this helps you out. Thank you! :)